OpenStudy (anonymous):

solve: y''-6y'+9y =3x^2

OpenStudy (anonymous):

are there boundary conditions? y(0) and y'(0)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the general solution is the sum of the homogeneous solution + particular solution. homogeneous: solution of y'' - 6y' +9y = 0 particular: solution to y'' - 6y' + 9y = 3x^2 I'll get to it very quickly but do you know how to find the solution to either one?

OpenStudy (anonymous):

I can find the first solution where y'' - 6y' +9y = 0, but i am getting confused with the particular solution

OpenStudy (anonymous):

for y'' - 6y' + 9y = f(x) the "form" of the particular solution we choose depends on what f(x) is. Since f(x) in this case is a 2nd order polynomial, the solution will have the form:\[y_p(x) = ax^2 + bx + c\]. Our goal is to determine a, b and c. To find it we plug it into the equation: y''(x) - 6y'(x) + 9y(x) = 3x^2 Let's find these first:\[y'(x) = 2ax + b\]\[y''(x) = 2a\] Now plug:\[2a - 6(2ax + b) + 9(ax^2 + bx + c) = 3x^2\] Simplified to:\[2a - 12ax -6b + 9ax^2 +9bx + 9c = 3x^2\] This line tells us that:\[9ax^2 = 3x^2\]\[(9b - 12a)x = 0x\]\[2a - 6b + 9c = 0\] First line: 9a = 3 ---> a = 1/3 Second line: 9b - 12a = 0 ---> b = 4/9 Third Line: 2(1/3) - 6(4/9) + 9c = 0 ---> c = 2/9 So the particular solution is:\[y_p(x) = \frac{ 1 }{ 3 }x^2 + \frac{ 4 }{ 9 }x +\frac{ 2 }{ 9 }\] General Solution: \[y_h(x) + y_p(x) = c_1 e^{3t} + c_2 te^{3t} + \frac{ 1 }{ 3 }x^2 + \frac{ 4 }{ 9 }x +\frac{ 2 }{ 9 }\]

OpenStudy (anonymous):

ok i see how that works now, just one tiny question on how you got : y′(x)=2ax+b y′′(x)=2a - other than that this has been huge help with understanding 2nd order problems

OpenStudy (anonymous):

y(x) = ax^2 + bx + c y'(x) = 2ax + b y''(x) = 2a glad i could help ^_^

OpenStudy (anonymous):

right, just taking derivative. i should have seen that. thanks again